# MCbiF: Measuring Topological Autocorrelation in Multiscale Clusterings via 2-Parameter Persistent Homology

### v1

@@ 4.2 MCbiF as a Higher-Order Sankey Diagram
Recall the definition of the Sankey diagram of the sequence of partitions introduced in Section 3.1, and its associated representation in terms of an $M$-layered graph with vertices $V_{m}$ at each layer representing the clusters of $\theta(t_{m})$, see Eq. (1). Let us define the disjoint union $A(\ell,m):=V_{\ell}\uplus...\uplus V_{m}$, $1\leq\ell\leq m$, which assigns an index to each cluster in $\theta(t)$ for $t\in[t_{\ell},t_{m}]$. Furthermore, recall that $\theta(t)_{i}$ denotes the $i$-th cluster $C_{i}$ of $\theta(t)$. The nerve-based MCbiF can then be defined as follows.
@@
The nerve-based MCbiF $\tilde{\mathcal{M}}$ is a 1-critical bifiltration with simplices representing clusters and their intersections, in contrast to the original MCbiF $\mathcal{M}$ (Eq (5)) in which the simplices represent elements in $X$ and their equivalence relations. Despite these different perspectives, Proposition 40 in the appendix shows that $\tilde{\mathcal{M}}$ and $\mathcal{M}$ lead to the same MPH and can be considered as equivalent. The proof of this equivalence follows from an extension of results by Schindler and Barahona [2025].
However, the dimensionality of $\mathcal{M}$ and $\tilde{\mathcal{M}}$ can differ, as shown in the following proposition.
@@
  1. (i)
    $\dim K^{s,t}=\max_{s\leq r\leq t}\max_{c\in\theta(r)}|C|-1$, $\forall t_{1}\leq s\leq t$.
  2. (ii)
    $\dim\tilde{K}^{t_{m},t_{m+n}}=n,\;\forall 1\leq m\leq M,0\leq n\leq M-m$.
See proof in Appendix A.2.2. The nerve-based MCbiF is therefore computationally advantageous when $M<\max_{t_{1}\leq t}\max_{C\in\theta(t)}|C|-1$, making it preferable in applications where the clusters are large.
The nerve-based MCbiF $\tilde{\mathcal{M}}$ can be interpreted as a higher-order extension of the Sankey diagram $S(\theta)$ (Eq. 1) in the sense that $\tilde{\mathcal{M}}$ not only records pairwise intersections between clusters in consecutive partitions of $\theta$, like $S(\theta)$), but also takes into account higher-order intersections between clusters in sub-sequences of $\theta$. More formally, we can state the following proposition that relates $S(\theta)$ to the nerve-based MCbiF.
@@
The Sankey diagram graph $S(\theta)$ is a strict 1-dimensional subcomplex of $\tilde{K}:=\tilde{K}^{t_{1},t_{M}}$. In particular, $V_{m}=K^{t_{m},t_{m}}$ and $E_{m}=K^{t_{m},t_{m+1}}$, $\forall m=1,\dots,M-1$. Hence, we can retrieve $S(\theta)$ from the zigzag filtration
@@
A proof can be found in Appendix A.2.2. For details on zigzag persistence, see Carlsson and de Silva [2010] and Appendix D.
Furthermore, the 0- and 1-conflicts that can arise in a single layer $E_{m}$ of the Sankey diagram can be fully characterised as follows.
@@
  1. (i)
    There is a 0-conflict in $t_{m},t_{m+1}]$ iff $\exists u\in V_{m}$ and $v\in V_{m+1}$ with $\mathrm{deg}(u)\geq 2$ and $\mathrm{deg}(v)\geq 2$, where $\mathrm{deg}$ denotes the node degree in the bipartite graph $(V_{m}\uplus V_{m+1},E_{m})$ associated with the Sankey diagram.
  2. (ii)
    There is a triangle 0-conflict in $t_{m},t_{m+1}]$ iff there is a path of length at least 3 in $E_{m}$.
  3. (iii)
    There is 1-conflict in $t_{m},t_{m+1}]$ iff there is an (even) cycle in $E_{m}$.
See Appendix A.2.2 for the proof and Fig. 3 for an illustration.
Figure 3: Relationship between different types of conflicts and the crossings in a single-layer Sankey diagram.
Importantly, a cycle in $E_{m}$ leads to a crossing in $E_{m}$ that cannot be undone, see Fig. 3d for an illustration. Hence, Proposition 28 (iii) implies that the sum of the elements of the superdiagonal of $\mathrm{HF}_{1}$ provides a lower bound for the minimal crossing number of the Sankey diagram, $\overline{\kappa}_{\theta}$ (defined in Eq. 2). We state this as the following corollary.
@@
Note that 1-conflicts that arise across multiple partitions in the sequence (i.e., across multiple layers) do not necessarily lead to crossings. See Fig. 1, where the chosen ordering of the elements does not lead to a crossing in the layout of the Sankey diagram despite the presence of a 1-conflict. However, we hypothesise that the full $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ feature maps capture more complicated crossings that arise in the Sankey layout across many layers. This insight is exploited in our computational tasks below.
@@ 5 Mathematical Links of MCbiF to Other Methods
We now present some mathematical connections of MCbiF to the related methods introduced in Section 2.
@@
Given a sequence of partitions $\theta$ with $\theta(t_{1}=0)=\hat{0}$ and $\theta(t_{M})=\hat{1}$, let us define the matrix of first-merge times conditioned on the starting scale $s$:
@@
Clearly, when $s=0$, this recovers the standard matrix of first-merge times $D_{\theta}:=D_{\theta,0}$ discussed in Section 2. If $\theta$ is hierarchical, i.e., an agglomerative dendrogram, then $D_{\theta}$ is an ultrametric, i.e., it fulfils the strong triangle inequality: $D_{\theta}(x,z)\leq\max\left(D_{\theta}(x,y),D_{\theta}(x,z)\right)$ $\forall x,y,z\in X$.
Corollary 16 states that the number of branches in the agglomerative dendrogram at level $t$, which is given by $|\theta(t)|$, is equal to $\mathrm{HF}_{0}(s,t)$ for any $s\leq t$. Hence, $\mathrm{HF}_{0}(s,t)$ contains the same information as the ultrametric in the hierarchical case, see also Schindler and Barahona [2025] and Proposition 32 below.
If, on the other hand, $\theta$ is non-hierarchical, triangle 0-conflicts can lead to violations of the (strong) triangle inequality:
@@
See Appendix A.3 for a proof and Fig. 8a for an illustration. Proposition 31 shows that $\bar{c}_{0}(\theta)$ measures how much the ultrametric property of $D_{\theta}$ is violated.
Recall that $D_{\theta,s}$ is a dissimilarity measure that can be used to define a filtration [Chazal et al., 2014]. Next, we show that the 0-dimensional MPH of MCbiF corresponds to the 0-dimensional MPH of a Rips-based bifiltration constructed from $D_{\theta,s}$.
@@
A proof is presented in Appendix A.3 and follows from an argument in Schindler and Barahona [2025]. In the hierarchical case, the 1-dimensional MPH of the Merge-Rips bifiltration is trivial because $D_{\theta}$ fulfils the strong triangle inequality, and is thus equivalent to the MCbiF, whose 1-dimensional MPH is also trivial in the hierarchical case, see Proposition 19.
Conditional Entropy.
The conditional entropy (CE) is only defined for pairs of partitions. CE is defined as the expected Shannon information of the conditional probability of a partition $\theta(t)=\{C_{1},\dots,C_{n}\}$ given $\theta(s)=\{C_{1}^{\prime},\dots,C_{m}^{\prime}\}$:
$P_{t|s}[i|j]=\frac{|C_{i}\cap C_{j}^{\prime}|}{|C_{j}^{\prime}|}.$
For the special case of $M=2$ (i.e., only two partitions in the sequence $\theta$), it can be shown that $\mathrm{HF}_{0}(t_{1},t_{2})$ follows directly from the spectral properties of the matrix $P_{t_{2}|t_{1}}P_{t_{2}|t_{1}}^{T}$ interpreted as an undirected graph, as shown by the following proposition.
Proposition 33.
$\mathrm{HF}_{0}(t_{1},t_{2})=\dim(\ker L)$, where $L=\mathrm{diag}(P_{t_{2}|t_{1}}\boldsymbol{1})-P_{t_{2}|t_{1}}P_{t_{2}|t_{1}}^{T}$ is a weighted Laplacian.
The proof can be found in Appendix A.3.
Remark 34.
Note that $P_{t|s}$ only encodes the pairwise relationship between clusters, and does not capture higher-order cluster inconsistencies. In particular, CE cannot detect 1-conflicts arising across more than two scales, as seen in Example 44 in Appendix B.
@@ 6.1 Regression Task: Minimal Crossing Number of Sankey Layout
In our first experiment, we consider a task of relevance in computer graphics and data visualisation: the minimisation of the crossing number of Sankey diagram layouts [Zarate et al., 2018, Li et al., 2025]. This minimisation is NP-complete, and here we use our MCbiF topological feature maps to predict the minimal crossing number $\overline{\kappa}_{\theta}$ of the Sankey diagram $S(\theta)$ of a given sequence of partitions $\theta$ (see Section 3.1).
Data
We test our measures on synthetic datasets generated by sampling randomly from the space of coarse-graining sequences of partitions.
Definition 35 (Space of coarse-graining sequences of partitions).
The space of coarse-graining sequences of partitions, denoted $\Pi_{N}^{M}$, is defined as the set of coarse-graining sequences $\theta:[0,\infty)\rightarrow\Pi_{X}$ with $|X|=N$ and $M$ change points $t_{m}=0,\ldots,M-1$, such that $|\theta(s)|\geq|\theta(t)|,\,\forall s\leq t$, which start with the partition of singletons $\theta(t_{1}=0)=\hat{0}$ and end with the full set $\theta(t_{M}=M-1)=\hat{1}$.
@@
Setting $M=20$, we generate two datasets of 20,000 random samples $\theta\in\Pi_{N}^{M}$ for $N=5$ and $N=10$. For each of the generated $\theta$, we compute three feature maps: the information-based pairwise conditional entropy matrix $\mathrm{CE}$ [Meilă, 2003] (see Eq. 15) and our MCbiF Hilbert functions ($\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$). In addition, as a baseline feature map, we also consider the (non-unique) raw label encoding of $\theta$ given by the $N\times M$ matrix whose $m$-th column contains the labels of the clusters in $\theta(t_{m})$ assigned to the elements in $X$. As our prediction target, we take $y=\overline{\kappa}_{\theta}$ (Eq. 3), the minimal crossing number of the layout of the Sankey diagram, which we computed with the OmicsSankey algorithm [Li et al., 2025]. See Section 3.1 for details. We expect that predicting $y$ will be harder for $N=10$ because the increased complexity of $\Pi_{M}^{N}$ allows for more complicated crossings in the Sankey diagram.
Table 1: Regression task. Test $R^{2}$ score of LR, CNN and MLP models trained on different features for $N=5$ and $N=10$. See Appendix F.1 for train $R^{2}$ scores.
NN Method Raw label encoding $\mathrm{HF}_{0}$ $\mathrm{HF}_{1}$ $\mathrm{HF}_{0}$ & $\mathrm{HF}_{1}$ $\mathrm{CE}$
5 LR 0.001 0.147 0.486 0.539 0.392
CNN -0.006 0.155 0.504 0.544 0.492
MLP -0.002 0.150 0.491 0.541 0.409
10 LR -0.012 0.214 0.448 0.516 0.457
CNN 0.000 0.211 0.448 0.507 0.454
MLP 0.000 0.212 0.450 0.514 0.458
Results
@@
As a preliminary assessment, we first compute the Pearson correlation, $r$, between the crossing number $y$ and the three measures under investigation: the information-theoretical measure $\mathrm{CI}$ (i.e., the average $\mathrm{CE}$, see Eq. (16)) and the MCbiF topological average measures $\bar{c}_{0}$ and $\bar{c}_{1}$. The correlation between $\mathrm{CI}$ and $y$ is low ($r=0.20$ for $N=5$ and $r=0.06$ for $N=10$), higher for $\bar{c}_{0}$, and highest for $\bar{c}_{1}$ ($r=0.47$ for $N=5,10$) (see Fig. 11 in Appendix F.1). This is consistent with our theoretical results in Section 4.2, which show the relation between the crossing number and $\mathrm{HF}_{1}$ (see, e.g., Corollary 29).
@@
We then proceed to the regression task of predicting $\overline{\kappa}_{\theta}$. We split each dataset into training (64%), validation (16%) and test (20%). For each feature map (or their combinations), we train three different models: linear regression (LR), multilayer perceptron (MLP), and convolutional neural network (CNN).We use the mean-squared error (MSE) as our loss function and employ the validation set for hyperparameter tuning. See Appendix F.1 for details of all the models. We then evaluate the model performance on the unseen test data using the coefficient of determination ($R^{2}$).
We find that the raw label encoding of $\theta$ does not improve upon the mean prediction ($R^{2}\approx 0$) and that MCbiF feature maps outperform the information-based feature map (Table 1). In particular, the combined $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ features lead to a significantly better model performance than $\mathrm{CE}$ ($p<0.0001$, $t$-test on the residuals). Furthermore, $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ yield $R^{2}=0.544$ for $N=5$ and $R^{2}=0.516$ for $N=10$ whereas $\mathrm{CE}$ only achieves $R^{2}=0.492$ and $R^{2}=0.458$, respectively. The strong performance of the simple LR model demonstrates the interpretability of the MCbiF features, important for explainable AI (XAI) [Adadi and Berrada, 2018].

### v2

@@ 3.2 MCbiF as a Higher-Order Sankey Diagram
Consider the definition of the Sankey diagram of the sequence of partitions provided in Section E.1, and its associated representation as an $M$-layered graph with vertices $V_{m}$ at each layer representing the clusters of $\theta(t_{m})$, see Eq. 9. Let us define the disjoint union $A(\ell,m):=V_{\ell}\uplus...\uplus V_{m}$, $1\leq\ell\leq m$, which assigns an index to each cluster in $\theta(t)$ for $t\in[t_{\ell},t_{m}]$. Furthermore, recall that $\theta(t)_{i}$ denotes the $i$-th cluster $C_{i}$ of $\theta(t)$. The nerve-based MCbiF can then be defined as follows.
@@
The nerve-based MCbiF $\tilde{\mathcal{M}}$ is a 1-critical bifiltration with simplices representing clusters and their intersections, in contrast to the original MCbiF $\mathcal{M}$ (Eq. 1) in which simplices represent elements in $X$ and their equivalence relations. Despite these different perspectives, Proposition 34 in Appendix B.2.2 shows that $\tilde{\mathcal{M}}$ and $\mathcal{M}$ lead to the same MPH and can be considered as equivalent. However, the dimensionality of $\mathcal{M}$ and $\tilde{\mathcal{M}}$ can differ, as shown in the following proposition.
Figure 2: Characterisation of conflicts in a single-layer Sankey diagram, following Proposition 18.
@@
(i) $\dim K^{s,t}=\max_{s\leq r\leq t}\max_{C\in\theta(r)}|C|-1$, $\forall t_{1}\leq s\leq t$. (ii) $\dim\tilde{K}^{t_{m},t_{m+n}}=n,\;\forall 1\leq m\leq M,0\leq n\leq M-m$.
See proof in Appendix B.2.2. The nerve-based MCbiF is therefore computationally advantageous when the number of scales $M$ is smaller than the size of the largest cluster, $M<\max_{C\in\theta([t_{1},\infty))}|C|$. The nerve-based $\tilde{\mathcal{M}}$ can be interpreted as a higher-order extension of the Sankey diagram $S(\theta)$ (Eq. 9). Yet, unlike $S(\theta)$, which only records pairwise intersections between clusters in consecutive partitions of $\theta$, $\tilde{\mathcal{M}}$ also accounts for higher-order intersections between clusters in sub-sequences of $\theta$.
@@
The Sankey diagram graph $S(\theta)$ is a strict 1-dimensional subcomplex of $\tilde{K}:=\tilde{K}^{t_{1},t_{M}}$. In particular, $V_{m}=\tilde{K}^{t_{m},t_{m}}$ and $E_{m}=\tilde{K}^{t_{m},t_{m+1}}$, $\forall m=1,\dots,M-1$. Hence, we can retrieve $S(\theta)$ from the zigzag filtration
@@
See proof in Appendix B.2.2. For details on zigzag persistence, see Carlsson & de Silva (2010) and Appendix E.4. Next, we characterise the 0- and 1-conflicts that can arise in a single layer $E_{m}$ of the Sankey diagram.
@@
(i) There is a 0-conflict in $t_{m},t_{m+1}]$ iff $\exists u\in V_{m}$ and $v\in V_{m+1}$ with $\mathrm{deg}(u)\geq 2$ and $\mathrm{deg}(v)\geq 2$, where $\mathrm{deg}$ denotes the node degree in the bipartite graph $(V_{m}\uplus V_{m+1},E_{m})$ associated with the Sankey diagram $S(\theta)$. (ii) There is a triangle 0-conflict in $t_{m},t_{m+1}]$ iff there is a path of length at least 3 in $E_{m}$. (iii) There is 1-conflict in $t_{m},t_{m+1}]$ iff there is a cycle in $E_{m}$.
See Appendix B.2.2 for a proof and Fig. 2 for an illustration. Importantly, a cycle in $E_{m}$ leads to a crossing in $E_{m}$ that cannot be undone ( Fig. 2d). Hence, Proposition 18 (iii) implies that the sum of the elements of the superdiagonal of $\mathrm{HF}_{1}$ provides a lower bound for the minimal crossing number of the Sankey diagram, $\overline{\kappa}_{\theta}$ defined in Eq. 10.
@@
Note that 1-conflicts that arise across multiple partitions in the sequence (i.e., across multiple layers of the Sankey diagram) do not necessarily lead to crossings. See Fig. 1 for a Sankey diagram with no crossing despite the presence of a 1-conflict. However, we hypothesise that the full $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ feature maps capture more complicated crossings that arise in the Sankey layout across many layers. This insight is exploited in our computational tasks below.
@@ 4 Mathematical Links of MCbiF to Other Methods
Given a sequence $\theta$ with $\theta(t_{1}=0)=\hat{0}$ and $\theta(t_{M})=\hat{1}$, let us define the matrix of first-merge times conditioned on the starting scale $s$:
@@
When $s=0$, this recovers the standard matrix of first-merge times $D_{\theta}:=D_{\theta,0}$ in Section 2. If $\theta$ is hierarchical, i.e., an agglomerative dendrogram, then $D_{\theta}$ is an ultrametric, as it fulfils the strong triangle inequality: $D_{\theta}(x,z)\leq\max\left(D_{\theta}(x,y),D_{\theta}(x,z)\right)$ $\forall x,y,z\in X$. From Corollary 9 we have that the number of branches in the agglomerative dendrogram at level $t$, which is given by $|\theta(t)|$, is equal to $\mathrm{HF}_{0}(s,t)$ for any $s\leq t$. Hence, $\mathrm{HF}_{0}(s,t)$ contains the same information as the ultrametric in the hierarchical case, see  Schindler & Barahona (2025) and Proposition 35. If, on the other hand, $\theta$ is non-hierarchical, triangle 0-conflicts can lead to violations of the (strong) triangle inequality.
@@
See Appendix B.3 for a proof and Fig. 7a for an illustration. Proposition 21 shows that $\bar{c}_{0}(\theta)$ measures how much the ultrametric property of $D_{\theta}$ is violated. Recall that $D_{\theta,s}$ is a dissimilarity measure that can be used to define a filtration (Chazal et al., 2014). We show in Proposition 35 in Appendix B.3 that the 0-dimensional MPH of MCbiF corresponds to the 0-dimensional MPH of the Merge-Rips bifiltration constructed from $D_{\theta,s}$. In the hierarchical case, the Merge-Rips bifiltration has trivial 1-dimensional MPH, as $D_{\theta}$ fulfils the strong triangle inequality, and is thus equivalent to the MCbiF, whose 1-dimensional MPH is also trivial in the hierarchical case, see Proposition 11.
Conditional Entropy.
The conditional entropy (CE) is only defined for pairs of partitions as the expected information of the conditional probability of $\theta(t)$ given $\theta(s)$, denoted $P_{t|s}$ (see Eq. 15 in Appendix F). For the case $M=2$ (i.e., only two partitions in the sequence $\theta$), $\mathrm{HF}_{0}(t_{1},t_{2})$ follows directly from the spectral properties of the matrix $P_{t_{2}|t_{1}}P_{t_{2}|t_{1}}^{T}$ interpreted as an undirected graph.
@@
$\mathrm{HF}_{0}(t_{1},t_{2})=\dim(\ker L)$ for graph Laplacian $L=\mathrm{diag}(P_{t_{2}|t_{1}}\bm{1})-P_{t_{2}|t_{1}}P_{t_{2}|t_{1}}^{T}$.
See Appendix B.3 for a proof. Note that $P_{t|s}$ only encodes the pairwise relationship between clusters, and does not capture higher-order cluster inconsistencies. In particular, CE cannot detect 1-conflicts arising across more than two scales, as seen in Example 39 in Appendix C.
@@ 5.1 Regression Task: Minimal Crossing Number of Sankey Layout
Our first experiment considers a task of relevance in computer graphics and data visualisation: minimising the crossing number of Sankey diagram layouts (Zarate et al., 2018; Li et al., 2025). This minimisation is NP-complete. Here we use our MCbiF feature maps to predict the minimal crossing number $\overline{\kappa}_{\theta}$ (Eq. 11) of the Sankey diagram $S(\theta)$ of a given sequence of partitions $\theta$ (see Section E.1).
@@
We build synthetic datasets by sampling randomly from $\Pi^{M}_{N}$, the space of coarse-graining sequences of partitions of $N$ elements with $M$ change points, see Definition 40 in Appendix D.1. Setting $M=20$, we generate two datasets of 20,000 random sequences $\theta\in\Pi_{N}^{M}$ for $N=5$ and $N=10$. For each $\theta$, we compute five feature maps: $\mathrm{CE}$, $\mathrm{ARI}$, $\mathrm{MOD}$ (see Section 2), and our MCbiF Hilbert functions ($\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$). To benchmark against representation learning, we also consider the (non-unique) raw label encoding of $\theta$ given by the $N\times M$ matrix whose $m$-th column contains the labels of the clusters in $\theta(t_{m})$ assigned to the elements in $X$, and the Sankey graphs $S(\theta)$ (Eq. 9). Our prediction target is $y=\overline{\kappa}_{\theta}$ (Eq. 11), the minimal crossing number of the layout of the Sankey diagram, as computed with the OmicsSankey algorithm (Li et al., 2025). See Section E.1 for details. We expect that predicting $y$ becomes harder for $N=10$ because the increased complexity of $\Pi_{M}^{N}$ allows for more complicated crossings in the Sankey diagram.
@@
As a preliminary assessment, we compute the Pearson correlation, $r$, between the crossing number $y$ and summary statistics of the five feature maps under investigation: the consensus indices $\overline{\mathrm{VI}}$, $\overline{\mathrm{ARI}}$ and $\overline{\mathrm{MOD}}$, and the MCbiF average measures $\bar{c}_{0}$ and $\bar{c}_{1}$. The correlation between the consensus indices and $y$ is low, higher for $\bar{c}_{0}$, and highest for $\bar{c}_{1}$ ($r=0.47$ for $N=5,10$) (see Fig. 10 in Appendix D.1). This is consistent with our theoretical results in Section 3.2, which show the relation between the crossing number and $\mathrm{HF}_{1}$ (see Corollary 19).
Table 1: Regression task. Test $R^{2}$ score of LR, CNN, MLP and GCN models trained on different feature sets for $N=5$ and $N=10$. See Appendix D.1 for confidence intervals and train $R^{2}$ scores.
Raw label Sankey
NN Method encoding graph $\mathrm{HF}_{0}$ $\mathrm{HF}_{1}$ $\mathrm{HF}_{0}$ & $\mathrm{HF}_{1}$ $\mathrm{CE}$ $\mathrm{ARI}$ $\mathrm{MOD}$
5 LR 0.078 - 0.147 0.486 0.539 0.392 0.166 0.413
CNN 0.267 - 0.155 0.504 0.544 0.492 0.422 0.354
MLP 0.104 - 0.150 0.491 0.541 0.409 0.214 0.351
GCN - 0.416 - - - - - -
10 LR 0.038 - 0.214 0.448 0.516 0.457 0.246 0.345
CNN 0.072 - 0.211 0.448 0.507 0.454 0.294 0.312
MLP 0.036 - 0.212 0.450 0.514 0.458 0.256 0.246
GCN - 0.229 - - - - - -
@@
For the regression task of predicting $\overline{\kappa}_{\theta}$, we split each dataset into training (64%), validation (16%) and test (20%). For each feature map (or their combinations), we train three models: linear regression (LR), multilayer perceptron (MLP), and convolutional neural network (CNN). We use the mean-squared error (MSE) as our loss function and employ the validation set for hyperparameter tuning. See Appendix D.1 for details. We then evaluate model performance on the unseen test data using the coefficient of determination ($R^{2}$). As representation learning from the raw data, we also train models (LR, CNN, MLP) on the raw label encodings, and a graph convolutional network (GCN) on the Sankey graph $S(\theta)$ (Eq. 9). Table 1 shows that models trained on MCbiF feature maps outperform all representation learning models; even the GCN trained on Sankey graphs ($R^{2}=0.416$ for $N=5$; $R^{2}=0.229$ for $N=10$) has significantly lower performance ($p<0.0001$, Wilcoxon test on residuals) than a simple LR model based on $\mathrm{HF}_{1}$ ($R^{2}=0.486$ for $N=5$; $R^{2}=0.185$ for $N=10$). This reflects the fact that the MCbiF is a higher-order extension of $S(\theta)$ that better captures the global properties of $\theta$ (see Proposition 17).
The MCbiF feature maps also outperform the baseline feature maps $\mathrm{CE}$, $\mathrm{ARI}$ and $\mathrm{MOD}$; the combined $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ features ($R^{2}=0.544$ for $N=5$; $R^{2}=0.516$ for $N=10$) have significantly better performance ($p<0.0001$, Wilcoxon test) than the best baseline features $\mathrm{CE}$ ($R^{2}=0.492$ for $N=5$; $R^{2}=0.458$ for $N=10$). Variability analysis (bootstrapped 95% confidence intervals) shows that models trained on the combined $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ features outperform the other models (see Table 3 in Appendix D.1). Note that the strong performance of simple LR models underscores the interpretability of the MCbiF features, important for explainable AI (Adadi & Berrada, 2018).